OpenStudy (anonymous):

Simplify ln((e^(3)/e^(x))

OpenStudy (anonymous):

Use the two rules: ln(a/b) = ln(a) - ln(b) ln(e^x) = x

OpenStudy (anonymous):

-x +3 ?

OpenStudy (anonymous):

I did ln (e^3 * e^-x) 3 + -x

OpenStudy (anonymous):

yea same thing

OpenStudy (anonymous):

But i added, you said ln(a/b) = ln(a) - ln(b) . So -X + 3 or x+3

OpenStudy (anonymous):

3 - x

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

ln(a/b) = ln(a) - ln(b) and ln(a*b) = ln(a) + ln(b)

OpenStudy (anonymous):

also I did ln (e^3 * e^-x) = ln(e^3/e^-x)

OpenStudy (anonymous):

sorry that was all jarbled

OpenStudy (anonymous):

ln(e^3 * e^-x) = ln(e^3/e^x)